Find the magnitude of the force between 1C and 2C that are 10 cm apart


1
Expert's answer
2022-01-14T09:46:02-0500

By Coulomb's law:


"|F|=\dfrac 1{4\pi\epsilon_0}\dfrac {q_1q_2}{R^2}=1.78·10^{12}\text{ N}."